OpenStudy (anonymous):

What is the value of x for the equation 2(4 + x) = 12? A)2 B)4 C)6 D)8 E)12

OpenStudy (anonymous):

4+x=6

OpenStudy (anonymous):

x=6-4 x=2

OpenStudy (anonymous):

Let's solve your equation step-by-step. 2(4+x)=12 Step 1: Simplify both sides of the equation. 2x+8=12 Step 2: Subtract 8 from both sides. 2x+8−8=12−8 2x=4 Step 3: Divide both sides by 2. 2x 2 = 4 2 x=2 Answer: x=2

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html is the website

OpenStudy (anonymous):

thank you my computer keeps freezing up. i hate this pellet

OpenStudy (anonymous):

lol your welcome

OpenStudy (anonymous):

http://www.acalculator.com/math-calculators.html